# Optimal Mechanisms for Demand Response: An Indifference Set Approach

### v1

@@ 5 Case Study: Pre-cooling in Phoenix, Arizona
Appendix A Proofs
@@ A.1 Proof of Proposition 1
$\widehat{\mathsf{R}}_{n}^{\mathsf{PM}}(p)=\frac{1}{n}\max\bigg{(}\sum\limits_{i=1}^{n}{\mathbb{I}}(a_{i}p_{1}\leq b_{i}p_{2})a_{i},\sum\limits_{i=1}^{n}{\mathbb{I}}(a_{i}p_{1}>b_{i}p_{2})b_{i}\bigg{)}\,.$
@@
$U_{n}^{+}=\min\limits_{0\leq z\leq 1}U_{n}^{+}(z)\,,\quad U_{n}^{+}(z)=\frac{1}{n}\max\bigg{(}\sum\limits_{i=1}^{n}{\mathbb{I}}\big{(}a_{i}z\leq b_{i}(1-z)\big{)}a_{i},\sum\limits_{i=1}^{n}{\mathbb{I}}\big{(}a_{i}z>b_{i}(1-z)\big{)}b_{i}\bigg{)}\,.$
@@
$U_{n}^{-}=\min\limits_{0\leq z\leq 1}U_{n}^{-}(z)\,,\quad U_{n}^{-}(z)=\frac{1}{n}\max\bigg{(}\sum\limits_{i=1}^{n}{\mathbb{I}}\big{(}-a_{i}z\leq b_{i}(-1+z)\big{)}a_{i},\sum\limits_{i=1}^{n}{\mathbb{I}}\big{(}-a_{i}z>b_{i}(-1+z)\big{)}b_{i}\bigg{)}$
@@
$\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\min\limits_{w_{i}\in[0,1]\,,\forall i\in[n]}\left\|\frac{1}{n}\sum\limits_{i=1}^{n}\begin{bmatrix}w_{i}a_{i}\\ (1-w_{i})b_{i}\end{bmatrix}\right\|_{\ell_{\infty}}\,.$
@@
$\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\min\limits_{w_{i}\in[0,1]\,,\forall i\in[n]}\max\limits_{\|r\|_{\ell_{1}}\leq 1}\frac{1}{n}\sum\limits_{i=1}^{n}\left(r_{1}w_{i}a_{i}+r_{2}(1-w_{i})b_{i}\right)\,.$
@@
$\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\max\limits_{\|r\|_{\ell_{1}}\leq 1}\min\limits_{w_{i}\in[0,1]\,,\forall i\in[n]}\frac{1}{n}\sum\limits_{i=1}^{n}\left(r_{1}w_{i}a_{i}+r_{2}(1-w_{i})b_{i}\right)\,.$
@@
$\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\max\limits_{\|r\|_{\ell_{1}}\leq 1}\frac{1}{n}\sum\limits_{i=1}^{n}\min\limits_{w_{i}\in[0,1]}\left(r_{1}w_{i}a_{i}+r_{2}(1-w_{i})b_{i}\right)\,.$
@@
$\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\max\limits_{\|r\|_{\ell_{1}}\leq 1}\frac{1}{n}\sum\limits_{i=1}^{n}\left(r_{1}a_{i}{\mathbb{I}}(r_{1}a_{i}\leq r_{2}b_{i})+r_{2}b_{i}{\mathbb{I}}(r_{1}a_{i}>r_{2}b_{i})\right)\,.$
@@
$\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}$ $=\max\limits_{0\leq z\leq 1}\frac{1}{n}\sum\limits_{i=1}^{n}\left((1-z)b_{i}{\mathbb{I}}\Big{(}(1-z)b_{i}<za_{i}\Big{)}+za_{i}{\mathbb{I}}\Big{(}(1-z)b_{i}\geq za_{i}\Big{)}\right)\,.$

### v2

@@ 5 Case Study: Pre-cooling in Phoenix, Arizona
6 Discussion
In this paper, we demonstrated the potential of price-based demand response mechanism to optimally leverage flexibility in consumer energy needs to achieve improved grid outcomes. In a departure from existing work (e.g., Adelman and Uçkun, 2019; Yang et al., 2012; Zhou et al., 2017, 2019), we modeled user flexibility using indifference sets that simply formalize high-level consumer preferences—and did not require consumers to reason about their full welfare function for energy use. Our analytic results under this the indifference set model were enabled by fundamental results on convex analysis for random sets (Aumann, 1965; Artstein and Vitale, 1975).
One interesting consequence of our indifference set model is that demand response becomes essentially a pure substitution phenomenon: Under our model, HEMS respond to price signals by shifting consumption across time periods rather than by curtailing overall consumption levels. We emphasize that other models of consumer price response do not result in such behavior. For example, Allcott (2011a) reports experimental results in a setting where consumers directly respond to time-varying energy prices (without using a HEMS or other technology device as an intermediary), and finds that people do not shift consumption across time: Instead, they simply reduced their consumption during peak hours. Whether or not consumers respond to time-varying prices by shifting consumption across time plays a key role in the overall effect of such programs (Jacobsen et al., 2020; Schittekatte et al., 2024). In policy discussions, it is important to distinguish between demand-response programs that assume consumers have the technological capacity to automatically respond to prices—as is done here or in, e.g., Adelman and Uçkun (2019) and Zhou et al. (2019)—versus programs where prices vary but consumers are simply expected to figure out how to respond to them on their own (Allcott, 2011a; Cosmo et al., 2014).
There are a number of questions still left open in this paper. Here, we take the consumer indifference sets, the total renewable production and the grid cost function as given. However, if our proposed pricing mechanism were adopted, all these elements may change endogenously. For example, if shifting energy use away from peak hours becomes both practical and profitable for consumers, they may be willing to increase the size of their indifference sets and thus provide further flexibility to the system; and shifts in consumer demand patterns may open the door to increased renewable adoption. It would be of considerable interest to study general-equilibrium implications of our proposed demand-response mechanism. Meanwhile, on a practical level, it’s clear that optimal time-varying prices may be different across days, depending on a number of factors such as weather, length of daylight and the day of the week. We are actively working on extensions of our learning algorithm that can accommodate and respond to such heterogeneity across days.
Appendix A Remaining Proofs
@@ A.1 Proof of Proposition 1
$\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{PM}}(p)=\frac{1}{n}\max\bigg{(}\sum\limits_{i=1}^{n}{\mathbb{I}}(a_{i}p_{1}\leq b_{i}p_{2})a_{i},\sum\limits_{i=1}^{n}{\mathbb{I}}(a_{i}p_{1}>b_{i}p_{2})b_{i}\bigg{)}\,.$
@@
$U_{n}^{+}=\min\limits_{0\leq z\leq 1}U_{n}^{+}(z)\,,\quad\frac{1}{n}U_{n}^{+}(z)=\frac{1}{n}\max\bigg{(}\sum\limits_{i=1}^{n}{\mathbb{I}}\big{(}a_{i}z\leq b_{i}(1-z)\big{)}a_{i},\sum\limits_{i=1}^{n}{\mathbb{I}}\big{(}a_{i}z>b_{i}(1-z)\big{)}b_{i}\bigg{)}\,.$
@@
$U_{n}^{-}=\min\limits_{0\leq z\leq 1}U_{n}^{-}(z)\,,\quad\frac{1}{n}U_{n}^{-}(z)=\frac{1}{n}\max\bigg{(}\sum\limits_{i=1}^{n}{\mathbb{I}}\big{(}-a_{i}z\leq b_{i}(-1+z)\big{)}a_{i},\sum\limits_{i=1}^{n}{\mathbb{I}}\big{(}-a_{i}z>b_{i}(-1+z)\big{)}b_{i}\bigg{)}$
@@
$\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\min\limits_{w_{i}\in[0,1]\,,\forall i\in[n]}\left\|\frac{1}{n}\sum\limits_{i=1}^{n}\begin{bmatrix}w_{i}a_{i}\\ (1-w_{i})b_{i}\end{bmatrix}\right\|_{\ell_{\infty}}\,.$
@@
$\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\min\limits_{w_{i}\in[0,1]\,,\forall i\in[n]}\max\limits_{\|r\|_{\ell_{1}}\leq 1}\frac{1}{n}\sum\limits_{i=1}^{n}\left(r_{1}w_{i}a_{i}+r_{2}(1-w_{i})b_{i}\right)\,.$
@@
$\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\max\limits_{\|r\|_{\ell_{1}}\leq 1}\min\limits_{w_{i}\in[0,1]\,,\forall i\in[n]}\frac{1}{n}\sum\limits_{i=1}^{n}\left(r_{1}w_{i}a_{i}+r_{2}(1-w_{i})b_{i}\right)\,.$
@@
$\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\max\limits_{\|r\|_{\ell_{1}}\leq 1}\frac{1}{n}\sum\limits_{i=1}^{n}\min\limits_{w_{i}\in[0,1]}\left(r_{1}w_{i}a_{i}+r_{2}(1-w_{i})b_{i}\right)\,.$
@@
$\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}=\max\limits_{\|r\|_{\ell_{1}}\leq 1}\frac{1}{n}\sum\limits_{i=1}^{n}\left(r_{1}a_{i}{\mathbb{I}}(r_{1}a_{i}\leq r_{2}b_{i})+r_{2}b_{i}{\mathbb{I}}(r_{1}a_{i}>r_{2}b_{i})\right)\,.$
@@
$\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}$ $=\max\limits_{0\leq z\leq 1}\frac{1}{n}\sum\limits_{i=1}^{n}\left((1-z)b_{i}{\mathbb{I}}\Big{(}(1-z)b_{i}<za_{i}\Big{)}+za_{i}{\mathbb{I}}\Big{(}(1-z)b_{i}\geq za_{i}\Big{)}\right)\,.$